### v1

# Communication Cost Reduction for Subgraph Counting under Local Differential Privacy via Hash Functions

Abstract
We suggest the use of hash functions to cut down the communication costs when counting subgraphs under edge local differential privacy. While various algorithms exist for computing graph statistics, including the count of subgraphs, under the edge local differential privacy, many suffer with high communication costs, making them less efficient for large graphs. Though data compression is a typical approach in differential privacy, its application in local differential privacy requires a form of compression that every node can reproduce. In our study, we introduce linear congruence hashing. With a sampling rate of $s$, our method can cut communication costs by a factor of $s^{2}$, albeit at the cost of increasing variance in the published graph statistic by a factor of $s$. The experimental results indicate that, when matched for communication costs, our method achieves a reduction in the $\ell_{2}$-error for triangle counts by up to 1000 times compared to the performance of leading algorithms.
I Introduction
Differential privacy [1, 2] has emerged as a benchmark for protecting user data. To meet this standard, it is necessary to obfuscate the publication results. This can be achieved by adding small noise [3] or altering the publication outcomes with small probability [4].
In differential privacy, it is typically assumed that there is a complete dataset available. After statistical analyses are done, obfuscation is applied to the results. However, there can be leaks of user information during the data collection or storage phases. To address these concerns, a variant of differential privacy, termed local differential privacy [5, 6], has been introduced. Here, rather than starting with a full dataset, each user is prompted to disguise their data before sharing it. As a result, the received dataset is not perfect. Numerous studies, like those referenced in [7, 8], have aimed to extract accurate statistical insights from these imperfect datasets. The local differential privacy has been practically employed by several companies to guarantee the privacy of their users’ information. Those companies include Apple, Microsoft, and Google [9, 10, 11].
Many studies on local differential privacy focus on tabular datasets, but there is also significant research dedicated to publishing graph statistics [12, 13]. For graph-based inputs, such as social networks, the prevalent privacy standard is edge local differential privacy [14]. In this framework, users are asked to share a disguised version of their adjacency vectors. These vectors, essentially bit sequences, signify mutual connections within the network. For instance, in a social network comprising $n$ users, user $v_{i}$ would provide their adjacency vector, denoted as $a_{i}=[a_{i,1},\dots,a_{i,n}]\in\{0,1\}^{n}$. Here, $a_{i,j}=1$ implies that users $v_{i}$ and $v_{j}$ are connected, while $a_{i,j}=0$ means there is no connection between them.
One widely adopted method for obfuscation is the randomized response [15, 16, 17]. In this approach, users are prompted to invert each bit in their adjacency vector, denoted as $a_{i,j}$, based on a specific probability. However, while the method is straightforward, its application to real-world social networks presents challenges. Typically, in practical social networks, an individual may have at most a few thousand friends, implying that the majority of $a_{i,j}$ values are zeroes. When employing the randomized response technique and flipping each bit of $a_{i,j}$ based on the designated probability, a significant number of zero bits get inverted to one. This phenomenon can distort the resulting graph statistics considerably [18, 19].
Beyond the publication of full graphs via randomized rounding, there is a growing interest in subgraph counting queries under local differential privacy. The challenge was first highlighted by [20], who proposed two key algorithms for $k$-star and triangle counting. Building on the $k$-star counting, [21] later introduced an unbiased and refined algorithm. For triangle counting, several noteworthy contributions include [22], which addressed the communication cost, [23], which delved into scenarios where users have access to a 2-hop graph view, and [24], which studies the lower bound of the additive error. Moving to the shuffle model, [25] gave algorithms for both triangle and 4-cycle counting. To conclude, [26] gave an approximation method for identifying the densest subgraph, rooted in locally adjustable algorithms.
Most of the works mentioned in the previous paragraph use the two-step mechanism [20, 22, 25]. In this mechanism, users are required to download the adjacency vectors of all other participants. They then compute graph statistics locally using their genuine adjacency vector and the obfuscated vectors from their peers. While this technique substantially boosts the accuracy for various graph statistics, such as the count of triangles, it comes with the drawback of demanding a vast amount of data download. Each user incurs a communication cost of $\Theta(n^{2})$. Given that the user count, $n$, can reach several billion in real-world scenarios, this download demand is not affordable for the majority.
Instead of using adjacency vectors, users might consider exchanging adjacency lists. This approach involves each user transmitting and receiving only a list of node pairs that are adjacent. If we assume that a user can have a limited, constant number of friends, then the bit-download requirement for the original graph becomes $\Theta(n\log n)$. However, given the surge in edge count to $\Theta(n^{2})$ due to the randomized response, the communication cost associated with downloading the obfuscated adjacency list shoots up to $\Theta(n^{2}\log n)$. This means that, post-obfuscation, the adjacency list fails to offer any communication efficiency benefits over the adjacency vector.
There are several techniques proposed to reduce the communication cost for the two-step mechanism such as the asymmetric randomized response [22], the degree-preserving randomized response [27], the degree-preserving exponential mechanism [28], or technique based on the compressive sensing [29]. However, we strongly believe that the communication cost could be further reduced both in theory and practice.
I-A Our Contributions
To mitigate the communication overhead, we suggest employing a compression technique on the adjacency vector or list prior to its transmission to the central server.
The concept of database compression is not novel within the realm of differential privacy. Numerous mechanisms have been developed to enhance the precision of statistics derived from tabular data through database compression using sampling. One of the most prominent among these is the SmallDB algorithm [30]. There have also been initiatives that harness the Kronecker graph to refine the accuracy of published graph statistics [31].
On the contrary, implementing database compression for the two-step mechanism under edge local differential privacy poses greater challenges. Imagine we compress the adjacency list $[a_{i,1},\dots,a_{i,n}]^{t}$ down to $[a_{i,c_{1}},\dots,a_{i,c_{m}}]^{t}$, with $m\ll n$, via sampling, and then disseminate this compressed list to all users. When other users utilize this compressed list for local calculations in the second step, they must be able to replicate the sampling outcome. We therefore need a function which can be replicated with small communication cost. This requirement inspired our proposition to utilize linear congruence hashing [32, 33] for the sampling process.
Using linear congruence hashing, we can evenly partition the node set $\{v_{1},\dots,v_{n}\}$ into $m$ subsets. Let us denote these subsets as $S_{1},\dots,S_{m}$, each having a size of $s$. The index $c_{i}$ is derived from $S_{i}$ with a uniform probability of $1/s$. Since all data is transmitted from the users with a probability of $1/s$, we can leverage the theorem on amplification by sub-sampling [34] to diminish the bit-flipping probability in the randomized response mechanism by a factor of $1/s$. Consequently, the count of non-zero entries in the bit vector $[a_{i,c_{1}},\dots,a_{i,c_{m}}]^{t}$ is roughly $m/s$, which is approximately equal to $n/s^{2}$. By employing a sampling rate of $s$, the communication overhead can be reduced by a factor of $s^{2}$.
While our publication’s variance might increase by up to a factor of $s$ for certain graph statistics, our experimental results demonstrate that, for a fixed download cost, our algorithm can reduce the $\ell_{2}$-error in triangle counts by a factor of 1000.
A valid inquiry might be whether we would obtain a comparable outcome by deterministically choosing the same disjoint sets $S_{1},\dots,S_{m}$ for every user $v_{i}$. While this approach could simplify the algorithm by eliminating the need for hash functions, our findings indicate that deterministic set selection leads to a significant variance in the estimations during the mechanism’s second step. Hence, using the linear congruence hashing is vital for our compression.
I-B Paper Organization
In Section II, we establish the fundamental notations of edge differential privacy and present the primitive mechanism that adheres to this privacy standard. Moving to Section III, we delve into the two-step mechanism, highlighting its communication costs in detail. Our approach, termed the group randomized response, is introduced in Section IV, which is then succeeded by a theoretical analysis in Section V. The methodology is applied to triangle counting in Section VI. We summarize and conclude our findings in Section VII.
II Preliminaries
II-A Hash functions
[…]
Each value in the hash space is referred to as a bin. Two keys, $\mathsf{x}$ and $\mathsf{y}$, are said to be in the same bin or to ‘collide’ if they have identical hash values. In other words, for a hash function $h$, a collision occurs if $h(\mathsf{x})=h(\mathsf{y})$. Hash functions are crafted to minimize the likelihood of collisions. Universal hashing, detailed in the subsequent definition, is a primary method employed in the design of these functions.
Definition 1 (Universal Hashing Scheme [36]).
A family of hash functions $\mathcal{H}$ is universal if, for all $\mathsf{x},\mathsf{y}\in[|0,n-1|]$ such that $\mathsf{x}\neq\mathsf{y}$, we have |$|\{h\in\mathcal{H}|h(\mathsf{x})=h(\mathsf{y})\}|\leq\frac{|\mathcal{H}|}{m}$.
It is shown in [36] that a scheme called linear congruence hashing [32, 33] is universal. The scheme can be defined as =$\mathcal{H}=\{h_{\mathsf{a},\mathsf{b}}|\mathsf{a}\in[|1,\mathsf{p}-1|],\mathsf{b}\in[|0,\mathsf{p}-1|]\}$ when $\mathsf{p}$ is a prime number greater than $n$ and $h_{\mathsf{a},\mathsf{b}}(\mathsf{x})=((\mathsf{a}\mathsf{x}+\mathsf{b})\text{mod }\mathsf{p})\text{ mod }m$. Notice that, for all $\mathsf{a,b}$ and $\mathsf{k}\in[|0,m-1|]$, we have $\left\lfloor\frac{n}{m}\right\rfloor\leq|\{\mathsf{x}:h_{\mathsf{a,b}}(\mathsf{x})=k\}|\leq\left\lceil\frac{n}{m}\right\rceil$. In other words, the number of key in a given bin is either $\left\lfloor\frac{n}{m}\right\rfloor$ or $\left\lceil\frac{n}{m}\right\rceil$.
II-B Edge Local Differential Privacy
Let the set of users be $\{v_{1},\dots,v_{n}\}$. Each user $v_{i}$ has own adjacency list $a_{i}=[a_{i,1},\dots,a_{i,n}]^{t}\in\{0,1\}^{n}$ where $a_{i,j}=1$ if $\{i,j\}\in E$ and $a_{i,j}=0$ otherwise. For any $a_{i},a_{i}^{\prime}\in\{0,1\}^{n}$, let $d(a_{i},a_{i}^{\prime})$ be the $\ell_{1}$-distance between the two vectors. The edge local differential privacy can be explained as in the following definition:
Definition 2 (Edge Local Differential Privacy [14]).
Let $\varepsilon>0$, let :$\mathcal{R}:\{0,1\}^{n}\rightarrow R$ be a randomized function, and let $\mathcal{A}:R^{n}\rightarrow\mathcal{S}$ be a function called aggregator. The mechanism $\mathcal{M}$ which outputs $\mathcal{A}(\mathcal{R}(a_{1}),\ldots,\mathcal{R}(a_{n}))$ is said to be $\varepsilon$-edge local differentially private if, for any $a_{i},a_{i}^{\prime}\in\{0,1\}^{n}$ such that $d(a_{i},a_{i}^{\prime})=1$ and any $\mathsf{R}\subset R$, we have $\mathbb{P}\left(\mathcal{R}(a_{i})\in\mathsf{R}\right)\leq e^{\varepsilon}\mathbb{P}\left(\mathcal{R}(a_{i}^{\prime})\in\mathsf{R}\right)$.
In this setting, $\varepsilon$ represents the degree of privacy protection and is termed the privacy budget. A lower privacy budget value indicates a stricter privacy limitation on published data. Each user possesses an adjacency list denoted by $a_{i}$. Before transmitting the result to the central server, they apply the randomized function $\mathcal{R}$. Subsequently, the central server aggregates this data using the aggregator $\mathcal{A}$. According to Definition 2, if a mechanism $\mathcal{M}$ is $\varepsilon$-edge differentially private, significant information of $a_{i}$ cannot be obtained from $\mathcal{R}(a_{i})$ once it is transferred and stored on the central server.
One of the most important properties of the edge local differential privacy is the composition theorem.
Theorem 1 (Composition Theorem [37]).
Let $\mathcal{M}_{1},\dots,\mathcal{M}_{p}$ be edge local differentially private mechanism with privacy budget $\varepsilon_{1},\dots,\varepsilon_{p}$. Then, the mechanism $\mathcal{M}_{p}\circ\cdots\circ\mathcal{M}_{1}$ is $(\varepsilon_{1}+\cdots+\varepsilon_{p})$-edge local differentially private.
II-C Basic Mechanisms
Numerous mechanisms have been suggested to meet Definition 2 [38, 39]. Among them, the edge local Laplacian mechanism [21] is an approach to offer privacy when we request each user to give real numbers to the aggregator. The Laplacian mechanism can be defined as in the following definition. It can be shown that the mechanism is $\varepsilon$-edge local differential private.
Definition 3 (Edge Local Laplacian Mechanism [21]).
For $f:\{0,1\}^{n}\rightarrow\mathbb{R}$, the global sensitivity of $f$ is $\Delta_{f}=\max\limits_{d(a,a^{\prime})=1}\lVert f(a)-f(a^{\prime})\rVert_{1}$. For $\varepsilon>0$, let $X$ be a random variable drawn from $\text{Lap}\left(\varepsilon/\Delta_{f}\right)$. The edge local Laplacian mechanism for $f$ with privacy budget $\varepsilon$ is the mechanism for which the randomized function $\mathcal{R}$ is defined as $f(a)+X$.
When we request users to send a vector of bits (such as the whole adjacency vector) to the aggregator, the edge local Laplacian mechanism is not the most suitable. Instead, a frequently used method is the randomized response [15, 16, 17], as detailed in the following definition.
Definition 4 (Randomized Response [15, 16, 17]).
For $\varepsilon>0$, the randomized response with privacy budget $\varepsilon$ is a mechanism which the randomized function :$\mathcal{R}:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$ is a function such that $\mathcal{R}(a_{i,1},\dots,a_{i,n})=(RR(a_{i,1}),\dots,RR(a_{i,n}))$ where
$\mathbb{P}\left(\text{RR}(a_{i,j})=1\right)=\begin{cases}\frac{e^{\varepsilon}}{1+e^{\varepsilon}}&\text{if }a_{i,j}=1\\ \frac{1}{1+e^{\varepsilon}}&\text{if }a_{i,j}=0.\end{cases}$ (1)
II-D Smooth Sensitivity
The traditional Laplacian Mechanism adjusts noise based on the global sensitivity, which is determined from the worst possible outcome for a selected function. It does not consider the specific adjacency list. Therefore, in certain situations, the global sensitivity might be much greater than the function’s changes due to slight update in the real adjacency list. Noticing this, [40] created a mechanism where the noise relates to the smooth sensitivity, which varies depending on the instance. But before we define smooth sensitivity, we first need to explain local sensitivity and sensitivity at a distance $k$.
Definition 5 (Local Sensitivity from Definition 1.4 of [40]).
For $f:\{0,1\}^{n}\rightarrow\mathbb{R}$, the local sensitivity of $f$ on $a\in\{0,1\}^{n}$ is $\text{LS}_{f}(a)=\max\limits_{a^{\prime}:d(a,a^{\prime})=1}\lVert f(a)-f(a^{\prime})\rVert_{1}$.
Simply put, the local sensitivity of $f$ on $\{0,1\}^{n}$ measures the largest change in $f$ due to a single edge alteration when starting with the adjacency list $a$. The sensitivity of $f$ at distance $k$ expands on this idea by looking at the changes resulting from a single alteration in the adjacency lists that are within $k$ modifications from $a$.
Definition 6 (Sensitivity at distance $k$ from Definition 3.1 of [40]).
For $f:\{0,1\}^{n}\rightarrow\mathbb{R}$, the sensitivity of $f$ at distance $k$ at $a$ is $A^{(k)}(a)=\max\limits_{a^{\prime}:d(a,a^{\prime})\leq k}\text{LS}_{f}(a^{\prime})$.
We are now ready to define the smooth sensitivity.
Definition 7 (Smooth Sensitivity from Definition 2.2 of [40]).
For $f$ a function and $\beta>0$, the $\beta$-smooth sensitivity of $f$ at $a$ is $S^{*}_{f,\beta}(a)=\max\limits_{k}e^{-\beta k}A^{(k)}(a)$
We are now ready to define the mechanism based on the smooth sensitivity. It is shown that the mechanism is $\varepsilon$-edge local differential private.
Definition 8 (Smooth Sensitivity mechanism from Lemma 2.5 of [40]).
Let $\gamma>1$ and $h(z)$ be a probability distribution proportional to $1/(1+|z|^{\gamma})$. We will denote $Z_{\gamma}$ the random variable drawn from $h$. For $\varepsilon>0$ and $\gamma>1$, the smooth sensitivity mechanism for $f$, denoted by $\mathcal{M}$, is a mechanism with the randomized function $\mathcal{R}(a)=f(a)+\frac{4\gamma}{\varepsilon}S^{*}_{f,\varepsilon/\gamma}(a)\cdot Z_{\gamma}$.
This result was later improved in [41] who proved that $\mathcal{R}(a)$ could be changed to $f(a)+\frac{2(\gamma-1)}{\varepsilon}S^{*}_{f,\varepsilon/2(\gamma-1)}(a)\cdot Z_{\gamma}$.
The mechanism operates effectively for all values of $\gamma>1$. Nonetheless, it is important to highlight that when $\gamma\leq 3$, we do not know a way to calculate the variance of the random variable $Z_{\gamma}$. Consequently, we will adopt $\gamma=4$ for this article, where the variance of $Z_{\gamma}$ is one.
III Previous Works on Subgraph Countings under Edge Local Differential Privacy
Let us suppose the the original graph is $G=(V,E)$ when $V=\{v_{1},\dots,v_{n}\}$ is a set of nodes or users and $E\subseteq\{\{u,v\}:u,v\in V\}$ is a set of edges or relationships. We can publish $G$ using the randomized response mechanism defined in the previous section. Given that $a^{\prime}_{i,j}=RR(a_{i,j})$ obtained from the randomized response, we can define a graph $G^{\prime}=(V,E^{\prime})$ with its adjacency matrix given by $A^{\prime}=(a^{\prime}_{i,j})_{1\leq i,j\leq n}$. While one can release the count of subgraphs by counting them in $G^{\prime}$, such a release is known to introduce a significant error [22].
III-A Two-Step Mechanism
To mitigate the error, Imola, Murakami, and Chaudhuri [20] introduced a two-step mechanism to release the count of triangles. A tuple of three nodes, represented as $(v_{i},v_{j},v_{k})$ where $i>j>k$ is a triangle if $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{j},v_{k}\}\in E$. We denote the set of triangles in $G$ by $\Delta_{G}$. Furthermore, for tuples in $\Delta_{G}$ where the initial element is $v_{i}$, we denote them as $\Delta_{i}$. We have that $|\Delta_{G}|=\sum\limits_{i}|\Delta_{i}|$.
The first step, represented by $\mathcal{M}_{1}$, involves releasing the obscured graph $A^{\prime}$. In the second step, represented by $\mathcal{M}_{2}$, every user $v_{i}$ provides an estimated value for $|\Delta_{i}|$ that is based on both the real adjacency vector $a_{i}$ and the distorted graph $G^{\prime}$. Let $\Delta^{\prime}_{i}=\{(v_{j},v_{k}):\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in G,\{v_{j},v_{k}\}\in G^{\prime},\text{ and }i>j>k\}$. Consider a function $f$ defined as $f(a_{i})=|\Delta^{\prime}_{i}|$. The randomized funtion of $\mathcal{M}_{2}$, represented as $\mathcal{R}_{2}$, aligns with what is used in the edge local Laplacian mechanism. Specifically, users are asked to add a Laplacian noise into $f(a_{i})$. The mechanism’s aggregator then sums the outcomes derived from these randomized functions.
The two-step mechanism can significantly increase the precision of the triangle counting. However, it has a large variance. Consider $i>i^{\prime}>j>k$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k}\}\in E$ and $\{v_{j},v_{k}\}\notin E$. If, by the randomized response, $\{v_{j},v_{k}\}\in E^{\prime}$, that would contribute two to the number of triangles. When the input graph has a large number of $C_{4}$, the variance becomes large. The authors of [22] propose a technique called four-cycle trick to reduce the variance. Define $\Delta_{i}^{\prime\prime}:=\{(v_{j},v_{k})\in\Delta_{i}^{\prime}:\{v_{i},v_{k}\}\in G^{\prime}\}$. Instead of having $f(a_{i})=|\Delta_{i}^{\prime}|$, they use $f(a_{i})=|\Delta_{i}^{\prime\prime}|$.
III-B Communication Cost
In many real-world social networks, each user is typically connected to a constant number of other users [42]. This means the number of connections (or ‘ones’ in the adjacency list $a_{i}$) is a lot fewer than $n$. So, instead of sending their entire adjacency list $a_{i}$ to the aggregator, user $v_{i}$ can just share the list of those directly connected to them. This approach reduces the communication cost from $n$ bits down to $\Theta(\log n)$.
However, when using randomized response mechanisms, every bit in $a_{i}$ has a constant probability of being flipped. The resulting adjacency list $a_{i}^{\prime}$ will have roughly $\Theta(n)$ ones. So, the communication-saving technique mentioned earlier will not work here.
The situation gets more difficult with the two-step mechanism discussed earlier. In this approach, each user needs to access the entire distorted graph $A^{\prime}$. The user $v_{i}$ cannot choose to access only the edge $\{v_{j},v_{k}\}$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E$ as that will reveal the existence of those two edges. This means they have to download the whole distorted graph, which requires about $\Theta(n^{2})$ bits. Such a cost becomes unrealistic for social networks with millions of nodes.
The four-cycle trick mentioned in the previous section can help reducing the communication cost. As a triangle $(v_{i},v_{j},v_{k})\in\Delta_{i}^{\prime\prime}$ only if $\{v_{i},v_{k}\}\in E^{\prime}$, user $v_{i}$ can calculate $|\Delta_{i}^{\prime\prime}|$ without accessing the information about whether $\{v_{j},v_{k}\}\in E^{\prime}$ or not when $\{v_{i},v_{k}\}\notin E^{\prime}$. Still, we think that the savings from the trick are not substantial enough to make the two-step method viable for big graphs.
IV Group Randomized Response
In this section, we introduce a mechanism designed for publishing graphs in a distributed environment under the edge local differential privacy. We have termed this approach the group randomized response (GRR). The mechanism is a four-step process:
  1. 1.
    Degree Sharing Step: In this step, each user share their degree protected by the Laplacian mechanism.
  2. 2.
    Grouping Step: In this step, nodes and edges are categorized into groups using the linear congruence hashing function.
  3. 3.
    Randomized Response Step: Here, obfuscation is applied solely to one randomly chosen representative from each group. Only edges corresponding to the representatives are sent to the aggregator.
  4. 4.
    Counting Step: Subgraphs are counted based on the information from the representatives.
Thanks to the sampling created by the grouping, this mechanism achieves a notably reduced communication cost. Moreover, the privacy budget amplification from the sampling [34] further decreases the probabilities of bit-flipping during the randomized response phase, thus further reduces the cost.
IV-A Degree Sharing Step
During this step, each user $v_{i}$ calculates the count of their connections, specifically |$|\{j<i:\{i,j\}\in E\}|$. They then share this count using the Laplacian mechanism with the privacy budget of $\varepsilon_{0}$. The shared count, denoted by $\tilde{d}_{i}$, will be used only for adjusting bias in the fourth step. If there are a lot of users ($n$ is large), any errors from this count are minor even if a large Laplacian noise is added. Thus, we can choose a relatively small value for $\varepsilon_{0}$.
IV-B Grouping Step
In this step, users do not publish any information. The aim here is to form random groups of edges that are of similar sizes. It is crucial that this randomization can be reproduced, and that the division is easily shareable among all participants. To achieve such a grouping that meets these criteria, we employ the linear congruence hashing, which is discussed in Section II.
[…]
Let $s$ be a parameter called sampling rate, $m$ be an integer such that $m=\lceil p/s\rceil$, and let $h_{i}(j)=h_{\theta_{i},\phi_{i}}(j)=((\theta_{i}\cdot j+\phi_{i})\text{ mod }p)\text{ mod }m$. Each user $v_{i}$ classifies the set of users $\{v_{1},\dots,v_{n}\}$ into $m$ disjoint groups, denoted by $S_{i,1},\dots,S_{i,m}$. The group $S_{i,t}$ can be defined as $\{v_{j}:h_{i}(j)=t\}$. By the property of the linear congruence hashing, we have that $s-1\leq|S_{i,t}|\leq s$. To make sure that all the sets have size $s$, we add a dummy element to all $S_{i,t}$ with size $s-1$. There is no edge between $v_{i}$ and the dummy node.
[…]
We then choose a representative for each group, $S_{i,t}$. Every member of the group has an equal probability of being selected, which is $1/s$. Let us call the chosen member from group $S_{i,t}$ as $v_{c_{t}}$. When $c_{t}$ is chosen from the group containing $k$, we will employ the notation $a_{i,h_{i}(k)}=a_{i,c_{t}}$ for convenience.
IV-C Randomized Response Step
Each user then sends the values $a_{i,c_{1}},\dots,a_{i,c_{m}}$ to the aggregator using the randomized response mechanism with a privacy budget of $\varepsilon^{\prime}=\ln(1+s(e^{\varepsilon_{1}}-1))$. It is noteworthy that this privacy budget, $\varepsilon^{\prime}$, is considerably greater than the $\varepsilon_{1}$ budget used in the traditional method. This results in a much smaller probability of flipping bits in the adjacency vector.
The obfuscated values sent from $v_{i}$ to the central system are denoted as $a^{\prime}_{i,c_{1}},\dots,a^{\prime}_{i,c_{m}}\in\{0,1\}$. The aggregator then releases these values, making them available for each user in the mechanism’s final step. In the earlier approach, users had to know the list of $v_{j}$ such that $a^{\prime}_{i,j}=1$. Now, they only need to know the list of $t$ such that $a^{\prime}_{i,c_{t}}=1.$ This significant reduces the number of bits users have to download from the central server. We will conduct the analysis on the number of bits which each user needs to download in the next section.
IV-D Counting Step
Recall that $d_{i}$ is the number of connections with nodes of smaller indexes of user $v_{i}$, which will be called low degree from here after. Let $p_{\text{p},i}:=\mathbb{P}\left(a_{j,h_{j}(k)}=1|a_{j,k}=1\right)$, i.e.
$p_{\text{p},i}=\frac{s-1}{s}\cdot\frac{d_{i}-1}{ms-1}+\frac{1}{s},$
and let $p^{\prime}_{\text{p},i}:=\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1|a_{j,k}=1\right)$, i.e.
$p^{\prime}_{\text{p},i}=\frac{e^{\varepsilon^{\prime}}}{1+e^{\varepsilon^{\prime}}}\cdot p_{\text{p},i}+\frac{1-p_{\text{p},i}}{1+e^{\varepsilon^{\prime}}}.$
Also, let $p_{\text{a},i}:=\mathbb{P}\left(a_{j,h_{j}(k)}=1|a_{j,k}=0\right)$, i.e.
$p_{\text{a},i}=\frac{s-1}{s}\cdot\frac{d_{i}}{ms-1},$
and let $p^{\prime}_{\text{a},i}:=\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1|a_{j,k}=0\right)$, i.e.
$p^{\prime}_{\text{a},i}=\frac{e^{\varepsilon^{\prime}}}{1+e^{\varepsilon^{\prime}}}\cdot p_{\text{a},i}+\frac{1-p_{\text{a},i}}{1+e^{\varepsilon^{\prime}}}.$
Define
$\omega_{i}$ ::=:= $\frac{1}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}$
$=$ $1/\left(\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}(p_{\text{p},i}-p_{\text{a},i})\right)$
$=$ $1/\left(\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}\left(\frac{1}{s}-\frac{1}{ms-1}\right)\right)$
$=$ $\frac{e^{\varepsilon^{\prime}}+1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}.$
Also, define
$\sigma_{i}$ ::=:= $\frac{p_{\text{a},i}^{\prime}}{p_{\text{p},i}^{\prime}-p_{\text{a},i}^{\prime}}$
$=$ $\left(\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}\cdot p_{\text{a},i}+\frac{1}{1+e^{\varepsilon^{\prime}}}\right)\cdot\frac{e^{\varepsilon^{\prime}}+1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}$
$=$ $p_{\text{a},i}\cdot\frac{ms^{2}-s}{ms-s-1}+\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}$
$=$ $\frac{s-1}{ms-s-1}d_{i}+\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}.$
Recall that $\tilde{d}_{i}$ is the low degree estimation of user $v_{i}$ published in the first step of our framework. Define
$\tilde{\sigma}_{i}:=\frac{s-1}{ms-s-1}\tilde{d}_{i}+\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}.$
We have that $\mathbb{E}[\tilde{\sigma}_{i}]=\sigma_{i}$. For all $i,t$, denote
$\tilde{a}_{i,c_{t}}:=\omega_{i}a_{i,c_{t}}^{\prime}-\tilde{\sigma}_{i}.$
In the previous work, subgraphs are counted based on the value of $a_{j,k}^{\prime}\in\{0,1\}$ obtained from the randomized mechanism. In GRR, we replace $a_{j,k}^{\prime}$ with $\tilde{a}_{j,c_{t}}$ when $t=h_{j}(k)$. Let $W_{i}:=\{(j,k):j<k<i\text{ and }\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E\}$. The estimated number of triangles (denoted by $|\Delta_{i}|$ was $\sum\limits_{(j,k)\in W_{i}}a^{\prime}_{j,k}$ in the previous work. In this work, the estimated number of triangle corresponding to $v_{i}$, denoted by $|\tilde{\Delta}_{i}|$, is $\sum\limits_{(j,k)\in W_{i}}\tilde{a}_{j,h_{j}(k)}$.
Similar to the previous method, each user calculate the number of subgraphs corresponding to their nodes. They then report the number using the Laplacian mechanism or the smooth sensitivity mechanism with the privacy budget $\varepsilon_{2}$. The estimated number of subgraphs is then equal to the sums of those reported from users.
V Theoretical Analysis of Our Mechanism
V-A Privacy
The privacy of our mechanism, discussed in the next theorem, is a direct result of the theorem on amplification by sub-sampling [34].
Theorem 2.
[…]
As the randomized response step has a privacy budget of $\varepsilon^{\prime}=\ln(1+s(e^{\epsilon_{1}}-1))$, the release of $a_{i,c_{1}},\dots,a_{i,c_{m}}$ in the first two steps has a privacy budget of $\ln(1+(e^{\varepsilon^{\prime}}-1)/s)=\ln(1+(1+s(e^{\varepsilon_{1}}-1)-1)/s)=\varepsilon_{1}$.
Since users do not release any information during the grouping step, the privacy budget of our four-step mechanism is then equal to $\varepsilon_{0}+\varepsilon_{1}+\varepsilon_{2}$ by the composition theorem.
V-B Expected Value of $\tilde{a}_{i,c_{t}}$
In the subsequent theorem, we discuss the expected value of $\tilde{a}_{i,c_{t}}$. The theorem reveals that when an estimate from the counting phase is a linear composition of $\tilde{a}_{i,c_{t}}$, it remains unbiased. The triangle count estimation in Section 4.3 is an example of these estimations.
Theorem 3.
For all $j,k\in\{1,\dots,n\}$, the expected value of $\tilde{a}_{j,h_{j}(k)}$ equals $a_{j,k}$.
Proof.
When $a_{j,k}=0$,
$\mathbb{E}\left[\tilde{a}_{j,h_{j}(k)}|a_{j,k}=0\right]$
$=$ $(\omega_{j}-\mathbb{E}[\tilde{\sigma}_{j}])\cdot p^{\prime}_{\text{a},j}-\mathbb{E}[\tilde{\sigma}_{j}]\cdot(1-p^{\prime}_{\text{a},j})$
$=$ $\omega_{j}p_{\text{a},j}^{\prime}-\sigma_{j}=0.$
When $a_{j,k}=1$, we have that:
$\mathbb{E}\left[\tilde{a}_{j,h_{j}(k)}|a_{j,k}=1\right]$
$=$ $(\omega_{j}-\mathbb{E}[\tilde{\sigma}_{j}])\cdot p^{\prime}_{\text{p},j}-\mathbb{E}[\tilde{\sigma}_{j}]\cdot(1-p^{\prime}_{\text{p},j})$
$=$ $\omega_{j}p^{\prime}_{\text{p},j}-\sigma_{j}=1.$
V-C Variance of $\tilde{a}_{i,c_{t}}$
The previous section demonstrated that our mechanism gives no bias for certain specific publications. In this subsection, we show that the variance of our mechanism is also quite minimal.
We use the following lemmas in the analysis.
Lemma 1.
For any $j,k$, $\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)=\frac{a_{j,k}+\sigma_{j}}{\omega_{j}}.$
Proof.
When $a_{j,k}=0$, we have that $\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)=p^{\prime}_{\text{a},j}=\frac{\sigma_{j}}{\omega_{j}}.$ When $a_{j,k}=1$, we have that $\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)=p^{\prime}_{\text{p},j}=\frac{1+\sigma_{j}}{\omega_{j}}.$
Lemma 2.
For all $i$ such that $d_{i}<m$, $\omega_{i}=O\left(\frac{s}{e^{\varepsilon_{1}}-1}\right)$ and $\sigma_{i}=O\left(\frac{1}{e^{\varepsilon_{1}}-1}\right)$.
Proof.
By $\varepsilon^{\prime}=\ln(1+s(e^{\varepsilon_{1}}-1))$, we have
$e^{\varepsilon^{\prime}}-1$ $=$ $s(e^{\varepsilon_{1}}-1)$
$\frac{1}{e^{\varepsilon^{\prime}}-1}$ $=$ $\frac{1}{s(e^{\varepsilon_{1}}-1)}.$
Hence,
$\omega_{i}$ $=$ $\frac{2+s(e^{\varepsilon_{1}}-1)}{s(e^{\varepsilon_{1}}-1)}\cdot\frac{ms^{2}-s}{ms-s-1}$
$=$ $O\left(\frac{s}{e^{\varepsilon_{1}}-1}\right).$
By the assumption that $d_{i}<m$, we have
$\sigma_{i}$$\leq$ $\frac{1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms^{2}-s}{ms-s-1}+1$
$=$ $\frac{1}{s(e^{\varepsilon_{1}}-1)}\cdot\frac{ms^{2}-s}{ms-s-1}+1$
$\leq$ $1+\frac{m}{m-1}\cdot\frac{1}{e^{\varepsilon_{1}}-1}=O\left(\frac{1}{e^{\varepsilon_{1}}-1}\right).$
We consider the variance of the variable $\tilde{a}_{j,h_{j}(k)}$ in the next theorem.
Theorem 4.
For all $j$ such that $d_{j}<m$, the variance of $\tilde{a}_{j,h_{j}(k)}$ is $O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}\right)$.
Proof.
We first consider the variance of the variable $a^{\prime}_{j,h_{j}(k)}$. Because the variable is binary, by Lemma 1, we obtain that
$\text{Var}\left(a^{\prime}_{j,h_{j}(k)}\right)$ $=$ $\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)\cdot\left(1-\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\right)\right)$
$=$ $\left(\frac{a_{j,k}+\sigma_{j}}{\omega_{j}}\right)\cdot\left(1-\frac{a_{j,k}+\sigma_{j}}{\omega_{j}}\right)$
$\leq$ $\frac{1+\sigma_{j}}{\omega_{j}}.$
Then, by Lemma 2,
$\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)$ $=$ $\text{Var}\left(\omega_{j}a^{\prime}_{j,h_{j}(k)}+\tilde{\sigma}_{i}\right)$
$=$ $\omega_{j}^{2}\text{Var}\left(a^{\prime}_{j,h_{j}(k)}\right)+\text{Var}\left(\tilde{\sigma}_{i}\right)$
$\leq$ $\omega_{j}(1+\sigma_{j})+\left(\frac{s-1}{ms-s-1}\right)^{2}\frac{2}{\varepsilon_{0}^{2}}$
$=$ $O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}\right).$
In the following subsection, we will elaborate that $\tilde{a}_{j,h_{j}(k)}$ can cut the communication cost by a factor of $s^{2}$ in contrast to using the direct randomized response results $a^{\prime}_{j,k}$. Given that the variance of $a^{\prime}_{j,k}$ is known to be $\Theta\left(\frac{1}{(e^{\epsilon_{1}}-1)^{2}}\right)$, the previous theorem suggests that, while we manage to decrease the communication cost by a factor of $s^{2}$, this is at the expense of amplifying the variance by a factor of $s$. In the upcoming section, we will demonstrate that, for a set communication cost, this mechanism yields a considerably reduced variance.
In situations where the goal is to sum up specific estimators, like in subgraph counting, a covariance emerges between these estimators. The size of this covariance is addressed in the subsequent theorem:
Theorem 5.
Assume $d_{i}<m$ for all $i$. The covariance of the estimators $\tilde{a}_{j,h_{j}(k)}$ and $\tilde{a}_{j^{\prime},h_{j^{\prime}}(k^{\prime})}$ is in $\mathcal{O}\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right)$ when $j=j^{\prime}$ and $k,k^{\prime}<j$. Otherwise, the covariance of the two estimators is zero.
Proof.
When $j\neq j^{\prime}$, the two estimators are sampled from different sets of variables. They are thus independent.
When $j=j^{\prime}$, the two estimators are identical when $k$ and $k^{\prime}$ are in the same bin, i.e. $h_{j}(k)=h_{j}(k^{\prime})$. This event, that we will denote $C^{(j)}_{k,k^{\prime}}$, happens with a probability of $\frac{s-1}{ms-1}$. As a result, for $k,k^{\prime}<j$:
$\text{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)$
=\ = $\omega_{j}^{2}\cdot\text{Cov}\left(a^{\prime}_{j,h_{j}(k)},a^{\prime}_{j,h_{j}(k^{\prime})}\right)$
=\ = $\omega_{j}^{2}\cdot\left(\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\right]-\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{\prime}_{j,h_{j}(k^{\prime})}\right]\right).$
Using those notations, we will consider the values of $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]$ and $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]$.
Recall from the previous section that $\mathbb{P}\left(a_{j,h_{j}(k)}=1\right)=\frac{s-1}{s}\cdot\frac{d_{j}}{ms-1}+a_{j,k}\left[\frac{1}{s}-\frac{1}{ms-1}\right]$. We have that:
$\frac{ms-1}{ms-2}\mathbb{P}\left(a_{j,h_{j}(k)}=1\right)$
$=$ $\frac{s-1}{s}\cdot\frac{d_{i}}{ms-2}+a_{j,k}\left[\frac{ms-1}{s(ms-2)}-\frac{1}{ms-2}\right]$
$\geq$ $\frac{s-1}{s}\cdot\frac{d_{i}-a_{j,k}}{ms-2}+\frac{a_{j,k}}{s}.$
For $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]$, let us denote $\mathbb{P}\left(a_{j,h_{j}(k)}=1\mid\overline{C^{(j)}_{k,k^{\prime}}}\right)$ as $p_{k}$. We obtain that:
$p_{k}$ $=$ $\frac{a_{j,k}}{s}+\frac{s-1}{s}\cdot\frac{d_{j}-a_{j,k}-a_{j,k^{\prime}}}{ms-2}$
$\leq$ $\frac{ms-1}{ms-2}\cdot\mathbb{P}\left(a_{j,h_{j}(k)}=1\right).$
Using those notations, we have:
$\mathbb{P}\left(a^{\prime}_{j,h_{j}(k)}=1\text{ and }a^{\prime}_{j,h_{j}(k^{\prime})}=1\mid\overline{C^{(j)}_{k,k^{\prime}}}\right)$
=\ = $\left(p_{k}\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}+\frac{1-p_{k}}{e^{\varepsilon^{\prime}}+1}\right)\left(p_{k^{\prime}}\frac{e^{\varepsilon^{\prime}}-1}{e^{\varepsilon^{\prime}}+1}+\frac{1-p_{k^{\prime}}}{e^{\varepsilon^{\prime}}+1}\right)$
[…]
In consequence, as long as $s$ is larger than 3, which will always be the case in practice, we have that:
$\mathbb{P}\left(\overline{C^{(j)}_{k,k^{\prime}}}\right)\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]\\ -\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{\prime}_{j,h_{j}(k^{\prime})}\right]\leq 0$
This simplifies the computation as we now have that:
$\text{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)\\ \leq\omega_{j}^{2}\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]\mathbb{P}\left(C^{(j)}_{k,k^{\prime}}\right)$
Now, let us denote $\mathbb{P}\left(a_{j,h_{j}(k)}=1\mid C^{(j)}_{k,k^{\prime}}\right)$ as $p^{\prime}_{k}$. We have that:
$p_{k}^{\prime}$ $=$ $\frac{a_{j,k}+a_{j,k^{\prime}}}{s}+\frac{s-2}{s}\cdot\frac{d_{j}-a_{j,k}-a_{j,k^{\prime}}}{ms-2}$
$\leq$ $\mathbb{P}\left(a_{j,h_{j}(k)}=1\right)+\mathbb{P}\left(a_{j,h_{j}(k^{\prime})}=1\right).$
Using the same argument as in the previous paragraph, we obtain that:
$\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}a^{\prime}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]$$\leq$ $\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]+\mathbb{E}\left[a^{\prime}_{j,h_{j}(k^{\prime})}\right].$
By Lemma 1 and the fact that $\mathbb{P}\left(C_{k,k^{\prime}}^{(j)}\right)=\frac{s-1}{ms-1}$, we get:
$\text{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)$$\leq$ $2\frac{s-1}{ms-1}\omega_{j}(1+\sigma_{j})$
$=$ $O\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right).$
It can be noticed that, when $s\ll n$, the covariance in the previous theorem is much smaller than the variance calculated in Theorem 4. We thus believe that the variance dominates almost all the loss we have found in the experiment.
V-D Communication Cost
The final aspect we will examine regarding the mechanism’s efficiency is the communication needed for graph aggregation and distribution to each user. We will particularly look at two metrics: the “download cost” (the average number of bits the server transmits to each user) and the “upload cost” (the average number of bits each user sends to the server). Our analysis will offer approximate values under the assumption that the graph is sparse, meaning the number of edges ($|E|$) is much less than the square of the number of vertices ($|V|^{2}$).
From our discussion in the previous section, it is evident that the majority of the upload cost arises from the randomized response step, while the download cost primarily stems from the counting step. Therefore, in this analysis, we will disregard the costs associated with the other steps.
Theorem 6.
The upload cost of user $v_{i}$ in our mechanism is $\mathcal{O}\left(\log{\frac{n}{s}}\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right)$.
Proof.
Recall that, in the randomized response step, each user $v_{i}$ submit the vector $[a^{\prime}_{i,c_{1}},\dots,a^{\prime}_{i,c_{m}}]\in\{0,1\}^{m}$ for $m\approx n/s$. To save the communication cost, the user can send the index set $\{c_{t}:a^{\prime}_{i,c_{t}}=1\}\subseteq\{1,\dots,m\}$. If $\ell$ elements of the vector are one, the cost is $\ell\cdot\log m\approx\ell\cdot\log n/s$.
Now, we discuss the expected value of $\ell$. As each $a^{\prime}_{i,j}=1$ is sampled in the vector with probability $1/s$ and $e^{\varepsilon^{\prime}}-1=s(e^{\varepsilon_{1}}-1)$, we have that
$\mathbb{E}[\ell]$ $=$ $\left(\frac{e^{\varepsilon^{\prime}}}{1+e^{\varepsilon^{\prime}}}d_{i}+\frac{1}{1+e^{\varepsilon^{\prime}}}(n-d_{i})\right)/s$
$\leq$ $\frac{d_{i}}{s}+\frac{1}{s^{2}(e^{\varepsilon_{1}}-1)}(n-d_{i})$
$=$ $O\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right).$
Next, we consider the download cost for our mechanism.
Corollary 1.
The download cost for all users is $\mathcal{O}\left(\log{\frac{n}{s}}\left(\frac{|E|}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right)$.
Proof.
As users have to download information from all others, the download cost is then:
$\sum_{i}\mathcal{O}\left(\log{\frac{n}{s}}\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right)$
$=$ $\mathcal{O}\left(\log{\frac{n}{s}}\left(\frac{|E|}{s}+\frac{n^{2}}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right).$
VI Application: Triangle Counting
While our GRR mechanism is adaptable for counting various subgraphs, the prevalent focus in existing literature, including [20, 22, 24], has been on triangle counting under local differential privacy. Therefore, in this section, we focus on triangle counting as well, enabling a direct comparison with these studies.
VI-A Sensitivity and Loss of Our Mechanism
Recall from the previous section that, in the counting step of our mechanism, we ask each user $v_{i}$ to estimate $|\Delta_{i}|$ when $\Delta_{i}=\{(v_{i},v_{j},v_{k}):i<j<k,\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{j},v_{k}\}\in E\}$. The number of triangles in the graph $G$, denoted by $|\Delta_{G}|$ is then the sum of all $|\Delta_{i}|$. The estimation of $|\Delta_{i}|$ in our mechanism is $\tilde{\Delta}_{i}:=\sum\limits_{j,k:\{i,j\},\{i,k\}\in E}\tilde{a}_{j,h_{j}(k)}$ added with the Laplacian noise when the variable $\tilde{a}_{j,h_{j}(k)}$ is defined in the previous section.
The magnitude of the Laplacian noise, dictated by the sensitivity of $\tilde{\Delta}_{i}$, is capped at $d_{i}\cdot\omega_{i}$, where $d_{i}$ is the degree of node $v_{i}$ and $\omega_{i}$ is the unbiased factor outlined in the previous section. If the degree $d_{i}$ is publicly known, we can directly set the sensitivity to $d_{i}\cdot\omega_{i}$. In scenarios where $d_{i}$ is not public, the sensitivity must be adjusted to the maximum degree, $d_{\text{max}}=\max_{i}d_{i}$. This maximum degree can be estimated as $\tilde{d}_{\text{max}}=C+\max_{i}\tilde{d}_{i}$, with $C$ being a positive integer and $\tilde{d}_{i}$ representing the degree announced in the degree sharing step. It is important to note, however, that this publication could potentially underestimate the actual maximum degree, leading to some bias in our publication. If a node $i$ has a degree $d_{i}$ that exceeds $\tilde{d}_{\text{max}}$, it is required to eliminate edges from its local data.
From now, let us suppose that $d_{i}$ is publicly known and $Lap_{i}$ is a noise sampled from the Laplacian distribution with parameter $\omega_{i}\cdot d_{i}/\varepsilon_{2}$. Consequently, the estimation of user $v_{i}$ is $\tilde{\Delta}_{i}+Lap_{i}$. In the preceding section, we established that $\tilde{\Delta}_{i}+Lap_{i}$ serves as an unbiased estimate of $|\Delta_{i}|$. The primary source of estimation loss arises from the variance of $\sum_{i}(\tilde{\Delta}_{i}+Lap_{i})$, which we consider in the subsequent theorem.
Theorem 7.
Let $S_{2}$, $C_{4}$, and $W_{4}$ are the number of two-star, four-cycle, and walks of length four in the graph $G$. The variance of our estimation $\sum_{i}(\tilde{\Delta}_{i}+Lap_{i})$ is $\mathcal{O}\left(\left(\frac{s}{(e^{\varepsilon_{1}}-1)\varepsilon_{2}}\right)^{2}\cdot S_{2}+\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\times(S_{2}+C_{4}+\frac{s}{n}\cdot W_{4})\right)$.
Proof.
The variance of $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}_{j,h_{j}(k)}+Lap_{i}\right)$ comes from four parts:
  1. 1.
    $\sum_{i}\text{Var}\left(Lap_{i}\right)$: By the parameter of the Laplacian noise, we know that the variance is $O\left(\sum_{i}\left(\frac{s\cdot d_{i}}{(e^{\varepsilon_{1}}-1)\varepsilon_{2}}\right)^{2}\right)=O\left(\left(\frac{s}{(e^{\varepsilon_{1}}-1)\varepsilon_{2}}\right)^{2}\cdot S_{2}\right)$;
  2. 2.
    $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)\right)$: We have from Theorem 4 that $\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)$ is $O\left(\frac{s}{(e^{e_{1}}-1)^{2}}\right)$. As there are $S_{2}$ terms in the summation, we have that the summation is no more than $O\left(\frac{s\cdot S_{2}}{(e^{e_{1}}-1)^{2}}\right)$.
  3. 3.
    The covariance resulting from the repeated counting of $\tilde{a}_{j,h_{j}(k)}$ in the summation: We observe that $\tilde{a}_{j,h_{j}(k)}$ might be counted twice if there exist $i,i^{\prime}>j>k$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{i^{\prime},k\}\in E$. From Theorem 4, the covariance for each instance of double counting is equal to $\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)=O\left(\frac{s}{(e^{e_{1}}-1)^{2}}\right)$, and the maximum frequency of such double counting is limited to no more than $C_{4}$.
  4. 4.
    The covariance between the terms $\tilde{a}_{j,h_{j}(k)}$ and $\tilde{a}_{j,h_{j}(k^{\prime})}$ arises in situations where both terms are included in the summation: This occurs when there are $i>j,k$ and $i^{\prime}>j,k^{\prime}$ such that the edge sets $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k^{\prime}}\}$ are part of $E$. These instances occur no more than $W_{4}$ times. According to Theorem 5, each occurrence contributes to the covariance with a value of $\frac{s^{2}}{n}\cdot\frac{1}{(e^{e_{1}}-1)^{2}}$.
VI-B Further Optimization for Triangle Counting: Central Server Sampling
We believe that the variance detailed in Theorem 7 is comparatively minimal. Nevertheless, specifically for triangle counting, this variance can be further reduced through a method we have termed “central server sampling.” The specifics of this technique will be outlined in this subsection.
VI-B1 Modification
Remember that through the GRR mechanism, the server acquires the set $\bar{\Delta}_{i}=\{t\mid a^{\prime}_{i,c_{t}}=1\}$ from user $v_{i}$ following the randomized response step. This set is then shared with all users at the start of the counting step. As identified in the prior subsection, this approach not only leads to significant communication costs but also contributes to a substantial covariance in triangle counting. To address these issues, we suggest generating subsets $\bar{\Delta}_{i}^{(1)},\dots,\bar{\Delta}_{i}^{(n)}$ from $\bar{\Delta}_{i}$ independently, and then distributing the subset $\bar{\Delta}_{i}^{(j)}$ to user $v_{j}$. An element $t$ in $\bar{\Delta}_{i}$ is independently selected for inclusion in $\bar{\Delta}_{i}^{(j)}$ with a constant probability, which we denote as $\mu_{c}$.
Let $a^{(i)}_{j,h_{j}(k)}\in\{0,1\}$ is a random variable indicating if $h_{j}(k)\in\bar{\Delta}_{j}^{(i)}$. The user $v_{i}$ believe that there is an edge $\{v_{j},v_{k}\}$ in the graph $G$ if $a^{(i)}_{j,h_{j}(k)}=1$. We have that $\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}\right]=\mu_{c}\cdot\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]$ for all $i,j,k$.
Recall the variables $\omega_{i}$ and $\tilde{\sigma}_{i}$ defined in the previous section. We define $\tilde{a}^{(i)}_{j,h_{j}(k)}=\frac{\omega_{i}}{\mu_{c}}a^{(i)}_{j,h_{j}(k)}+\tilde{\sigma}_{i}$.
Theorem 8.
The variable $\tilde{a}^{(i)}_{j,h_{j}(k)}$ is an unbiased estimation of $a_{j,k}$.
Proof.
By the definitions and Theorem 3, we have that
$\mathbb{E}[\tilde{a}^{(i)}_{j,h_{j}(k)}]$ $=$ $\frac{\omega_{i}}{\mu_{c}}\mathbb{E}[a^{(i)}_{j,h_{j}(k)}]+\mathbb{E}[\tilde{\sigma}_{i}]$
$=$ $\omega_{i}\mathbb{E}[a^{\prime}_{j,h_{j}(k)}]+\mathbb{E}[\tilde{\sigma}_{i}]$
$=$ $a_{j,k}.$
From the previous theorem, our unbiased estimation of the number of triangles under this modification is then $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}^{(i)}_{j,h_{j}(k)}+Lap_{i}\right)$ when $Lap_{i}$ is a noise sampled from the Laplacian distribution with parameter $d_{i}\cdot\omega_{i}/(\mu_{c}\cdot\varepsilon_{2})$.
VI-B2 Variance of the Modified Mechanism
We first consider the variance of $\tilde{a}^{(i)}_{j,h_{j}(k)}$ in the subsequent theorem:
Theorem 9.
[…]
Proof.
Because $a^{(i)}_{j,h_{j}(k)}$ is a binomial variable, we have from Lemma 1 that:
$\text{Var}\left(a^{(i)}_{j,h_{j}(k)}\right)$
$=$ $\mathbb{P}\left(a^{(i)}_{j,h_{j}(k)}=1\right)\cdot\left(1-\mathbb{P}\left(a^{(i)}_{j,h_{j}(k)}=1\right)\right)$
$=$ $\mu_{c}\cdot\mathbb{P}\left({a^{\prime}}_{j,h_{j}(k)}=1\right)\cdot\left(1-\mu_{c}\cdot\mathbb{P}\left(a^{(i)}_{j,h_{j}(k)}=1\right)\right)$
$\leq$ $\mu_{c}\cdot\mathbb{P}\left({a^{\prime}}_{j,h_{j}(k)}=1\right)$
$\leq$ $\mu_{c}\frac{1+\sigma_{j}}{\omega_{j}}.$
Then, we have from Lemma 2 that:
$\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$ $=$ $\frac{\omega_{j}^{2}}{\mu_{c}^{2}}\text{Var}\left(a^{(i)}_{j,h_{j}(k)}\right)+\text{Var}\left(\tilde{\sigma}_{i}\right)$
$\leq$ $\frac{1}{\mu_{c}}\frac{1+\sigma_{j}}{\omega_{j}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}$
$=$ $O\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}\right).$
At first glance, when contrasting this outcome with the findings in Theorem 4, it might appear that employing central server sampling amplifies the variance of the estimations, potentially diminishing their utility. However, it is important to recognize that for a communication cost reduction by $1/\mu_{c}$ times. Additionally, in the same way as the four-cycle trick in Section III, the central server sampling can reduce the covariance stated in Theorem 7. In particular, it significantly reduces the covariance for the graph with a lot of walks with length four (or graph with large number of $W_{4}$).
Theorem 10.
The covariance of the estimators $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j^{\prime},h_{j}^{\prime}(k^{\prime})}$ is $O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\right)$ when $j=j^{\prime}$, $i\neq i^{\prime}$ and $k=k^{\prime}$, $O\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right)$ when $j=j^{\prime}$, $i\neq i^{\prime}$ and $k,k^{\prime}<j$, and $O\left(\frac{s^{2}}{\mu_{c}n(e^{\varepsilon_{1}}-1)^{2}}\right)$ when $j=j^{\prime}$, $i=i^{\prime}$ and $k,k^{\prime}<j$. It is equal to zero in other cases.
Proof.
When $j\neq j^{\prime}$, the two estimators are sampled from different sets of variables. Furthermore, the server sampling step is done independently for each group. The estimators are thus independent.
When $j=j^{\prime}$, $i\neq i^{\prime}$ and $k,k^{\prime}<j$, the server sampling step is done independently for $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}$. Thus, by Theorem 5,
$\text{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}\right)$$\leq$ $\text{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)$
$=$ $O\left(\frac{s^{2}}{n(e^{\epsilon_{1}}-1)^{2}}\right).$
Similarly, when $j=j^{\prime}$, $i\neq i^{\prime}$ and $k=k^{\prime}$, we have from Theorem 4 that:
$\text{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i^{\prime})}_{j,h_{j}(k)}\right)\leq\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)=O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\right).$
When $j=j^{\prime}$, $i=i^{\prime}$ and $k,k^{\prime}<j$, we will conduct the analysis by separating the case when $h_{j}(k)=h_{j}(k^{\prime})$ and the case when $h_{j}(k)\neq h_{j}(k^{\prime})$. Recall that $C_{k,k^{\prime}}^{(j)}$ is the event when $h_{j}(k)=h_{j}(k^{\prime})$. We have from Theorem 9 that:
$\mathbb{E}\left[\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\mid C_{k,k^{\prime}}^{(j)}\right]$ $=$ $\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$
$=$ $O\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\right).$
In the case where $k$ and $k^{\prime}$ are not colliding, we straightforwardly obtain from the proof of Theorem 5 that $\mathbb{P}\left(\overline{C^{(j)}_{k,k^{\prime}}}\right)\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}a^{(i)}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]-\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{(i)}_{j,h_{j}(k^{\prime})}\right]\leq 0$. Hence, we can say that:
$\text{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\right)$
$\leq$ $\mathbb{P}\left(C^{(j)}_{k,k^{\prime}}\right)\mathbb{E}\left[\tilde{a}^{(i)}_{j,h_{j}(k)}\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]$
$=$ $O\left(\frac{s^{2}}{\mu_{c}n(e^{\varepsilon_{1}}-1)^{2}}\right).$
Using the previous two theorems, we will quantify the variance of the triangle counting algorithm performed using GRR and central server sampling. We describe in the next theorem this variance prior to the addition of the Laplacian noise in the final step of the mechanism.
Theorem 11.
Let $S_{2}$, $S_{3}$, $C_{4}$, and $W_{4}$ be the number of two-star, three-star, four-cycle, and the walk with length four in $G$. The variance of our estimator for triangle counting using the GRR and central server sampling technique, denoted by $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$, is $\mathcal{O}\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\left(S_{2}+\mu_{c}C_{4}+\frac{\mu_{c}s}{n}W_{4}+\frac{s}{n}S_{3}\right)\right)$.
Proof.
The variance of $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$ comes from four parts:
  1. 1.
    $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)\right)$: We have from Theorem 9 that $\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$ is $O\left(\frac{s}{\mu_{c}(e^{e_{1}}-1)^{2}}\right)$. As there are $S_{2}$ terms in the summation, we have that the summation is no more than $O\left(\frac{s\cdot S_{2}}{\mu_{c}(e^{e_{1}}-1)^{2}}\right)$.
  2. 2.
    The covariance resulting from the repeated counting of $\{j,k\}$ in the summation: This is from the covariance of $a_{j,h_{j}(k)}^{(i)}$ and $a_{j,h_{j}(k)}^{(i^{\prime})}$ for $i,i^{\prime}>j>k$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k}\}\in E$. By Theorem 10, the covariance of $a_{j,h_{j}(k)}^{(i)}$ and $a_{j,h_{j}(k)}^{(i^{\prime})}$ is $O\left(\frac{s}{(e^{e_{1}}-1)^{2}}\right)$, and the maximum frequency of such double counting is limited to no more than $C_{4}$.
  3. 3.
    The covariance between the terms $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}$ for $i\neq i^{\prime}$ where both terms are included in the summation: This occurs when there are $i>j,k$ and $i^{\prime}>j,k^{\prime}$ such that the edge sets $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k^{\prime}}\}$ are part of $E$. These instances occur no more than $W_{4}$ times. According to Theorem 5, each occurrence contributes to the covariance with a value of $\frac{s^{2}}{n}\cdot\frac{1}{(e^{e_{1}}-1)^{2}}$.
  4. 4.
    The covariance between the terms $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}$ where both terms are included in the summation: This occurs when there are $i>j,k$ such that the edge sets $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i},v_{k^{\prime}}\}$ are part of $E$. These instances occur no more than $S_{3}$ times. According to Theorem 5, each occurrence contributes to the covariance with a value of $\frac{s^{2}}{\mu_{c}n}\cdot\frac{1}{(e^{e_{1}}-1)^{2}}$.
Let us compare the variance derived from the Generalized Random Response (GRR) as outlined in Theorem 7 with that obtained from GRR combined with central server sampling, as detailed in Theorem 11, particularly for graphs with numerous walks of length four. In such scenarios, the term $W_{4}$ becomes the dominant factor in the variances of both methods. We find that the constant associated with this dominant term is identical in both variances. This observation suggests that while achieving a similar level of variance, we can also reduce the download cost by a factor of $1/\mu_{c}$.
VI-C Smooth Sensitivity for Triangle Counting
As mentioned in the preceding subsection, the clipping method can be employed to diminish the magnitude of Laplacian noise introduced in the final stage of our mechanism. However, this method of clipping incurs a certain degree of bias in our estimator. To remove this bias, we suggest adopting the smooth sensitivity approach, which is introduced in Section II.
We now calculate the sensitivity at distance $k$ (defined in Definition 6) for the triangle counting. First, let us consider the local sensitivity (defined in Definition 5) for the adjacency vector $a^{\prime}\in\{0,1\}^{n}$ (denoted by $LS_{f}(a^{\prime})$). As previously discussed, the maximum change in the count $\sum\limits_{j,k:a^{\prime}_{j}=a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}$ when one bit of $a^{\prime}$ is updated is no more than $UB_{\text{LS}}(a^{\prime})=\max\limits_{j}\left(\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}+\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{k,h_{k}(j)}\right)$.
When $a_{i}$ is the adjacent vector of user $i$ and $d_{i}$ is the number of ones in the vector, the number of ones in the vector $a$ with $|a-a_{i}|\leq k$ is not more than $UB_{\text{LS}}(a_{i})+k\cdot\omega_{i}$. Hence, the sensitivity at $k$ of the vector $a_{i}$, denoted by $A^{(k)}(a_{i})$ is no more than $O\left(UB_{\text{LS}}(a_{i})+\frac{k\cdot s}{e^{\varepsilon_{1}}+1}\right)$.
VI-D Experiments
Figure 1: Comparison on relative error between our algorithm and ARR for various download and upload costs on the Wikipedia Article Network graph with $\varepsilon=1$. In the legend, the abbreviation ’CSS’ denotes our Central Server Sampling method.
In this subsection, we aim to compare the algorithm we’ve outlined in earlier sections, which is fully accessible via our GitHub repository at https://github.com/Gericko/GroupRandomizedResponse, with the leading-edge algorithm named ARROne from [22]. It is worth noting that our implementation of ARROne differs slightly from the original version described in the cited paper, as we believe the original underestimates the sensitivity [43]. Specifically, in our version, the contribution of an edge $(v_{i},v_{j})$ to the count by user $v_{i}$ is defined as |$|\{k:a_{i,k}=1,\{v_{j},v_{k}\}\in M_{i},\text{ and }k<i\}|$ rather than |$|\{k:a_{i,k}=1,\{v_{j},v_{k}\}\in M_{i},\text{ and }j<k<i\}|$. Among the various algorithms discussed in [22], we focus on the one employing the 4-cycle trick, as it gives the highest performance.
Our experiments were performed on the Wikipedia Article Networks dataset [44, 45]. This dataset models Wikipedia users as nodes and their votes for admin promotions as edges. We selected this graph, which is based on real-world data, to assess the accuracy of our algorithm in practical scenarios. The Wikipedia dataset comprises 7,115 nodes and 100,762 edges. For experiments involving different graph sizes from the original, we created the extracted graph by randomly selecting a set of nodes of the desired size and then taking the subgraph induced by that set.
Our experiments focus on tracking the accuracy, quantified by relative error, as we modify various parameters. These parameters include the download cost, upload cost, size of the graph, and privacy budget.
The values presented for each parameter set represent the average derived from 10 separate simulation runs.
VI-D1 Error Analysis for Various Download Costs
The first experiment involves a comparison between the following three methods:
  1. 1.
    The leading edge method detailed in [22]
  2. 2.
    Our GRR mechanism enhanced by the central server sampling and the clipping method [^1] The clipping method can also be applied in the counting step of the Generalized Random Response (GRR). We have established a theorem that specifies a clipping factor, ensuring that an edge is only clipped with a minimal probability. Nonetheless, we have opted not to delve into the details of the clipping factor in this paper, as our experiments indicate that the clipping method is outperformed by the application of smooth sensitivity. We have incorporated the results of the GRR mechanism, augmented with the clipping method, to verify and demonstrate its superiority over the ARR method.
  3. 3.
    Our GRR mechanism enhanced by the central server sampling and the smooth sensitivity
Figure 1 illustrates that our algorithm outperforms the leading-edge algorithm across all download costs. When the download cost is minimal, the enhancement in relative error can be as substantial as a 1000-fold reduction.
The performance gap between our algorithm and the one presented in [22] widens as the communication cost decreases. This is attributed to the sampling rate $\mu^{*}$ dependency. To recap, $\mu^{*}$ represents the equivalent download cost reduction, equating to $\mu_{c}/s^{2}$ for GRR with central server sampling, and $\mu^{2}$ for ARROne. In the case of [22], the $S_{2}$ term scales with $1/\mu^{*}$ and the $C_{4}$ term with $1/\sqrt{\mu^{*}}$. For our algorithm, when setting $s=(1/\mu^{*})^{1/3}$ and $\mu_{c}=(\mu^{*})^{1/3}$, the $S_{2}$ term scales with $(\mu^{*})^{-2/3}$ and the $C_{4}$ term with $(\mu^{*})^{-1/3}$. This scaling results in our algorithm being more effective at lower communication costs compared to the approach in [22].
When the download cost reaches its maximum — a scenario that occurs when no sampling is implemented — our mechanism and the leading mechanism exhibit similar performance. This outcome is anticipated, as in this scenario, both approaches essentially align with each other and with the classical local differential privacy-based triangle counting algorithm, as described in [20].
The experimental findings also reveal that, across all the download cost budgets we considered, the smooth sensitivity method consistently outperforms the clipping method.
VI-D2 Error Analysis for Upload Costs
The second experiment involves a comparison between the following three methods:
  1. 1.
    The leading edge method detailed in [22]
  2. 2.
    Our GRR mechanism enhanced the clipping method
  3. 3.
    Our GRR mechanism enhanced the smooth sensitivity
In this experiment, we have opted not to implement the four-cycle trick or the central server sampling, as neither method contributes to the reduction of upload costs.
Figure 1 demonstrates that our approach has enhanced the relative errors in counting across all ranges of upload cost budgets. Notably, the smaller the budget, the more significant the improvement we observe. For the lowest upload cost budget in this experiment, we have achieved an enhancement in relative error by up to a factor of 1000. It is also evident that the smooth sensitivity method exhibits superior performance compared to the clipping method in this context.
We can notice from the results that the error of our GRR mechanism tends to plateau at higher upload costs. This phenomenon can be attributed to the behavior of the variance term associated with the publishing of the triangle count, which becomes significant for moderately sized graphs. Specifically, with minimal sampling, the number of edges in the obfuscated graph diminishes at a rate comparable to the increase of $\omega$. Consequently, the local sensitivity remains relatively constant. However, as the sampling extent increases, the edge count in the obfuscated graph becomes too sparse for the law of large numbers to be effective.
As the sampling increases, the sum $\sum\limits_{k:a^{\prime}_{k}=1}a^{(i)}_{j,h_{j}(k)}$ becomes a collection of random variables where the likelihood of each being 1 is considerably less than the inverse of the number of terms in the sum. Consequently, most often this sum will equal 0, with occasional instances of 1 or 2. However, the maximum value of this sum across all $j$ will almost invariably be 1 or 2. This leads to the upper bound on the local sensitivity, denoted as $UB_{\text{LS}}(a^{\prime})=\max\limits_{j}\left(\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}+\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{k,h_{k}(j)}\right)$, being equivalent to either $\omega_{j}$ or $2\omega_{j}$. Given that $\omega_{j}$ scales linearly with the sampling factor, it follows that for high levels of sampling, the noise added to the count also increases almost linearly with the sampling factor.
VI-D3 Error Analysis for Various Graph Sizes and Privacy Budgets
Figure 2: Comparison on relative error between our algorithm and the leading edge algorithm for various graph sizes generated from the Wikipedia Article Network graph with $\varepsilon=1$ and a download cost reduction of 1000. In the legend, the abbreviation ’CSS’ denotes our Central Server Sampling method.
Figure 3: Comparison on relative error between our algorithm and the leading edge algorithm for various privacy budgets on the Wikipedia Article Network with a download cost reduction of 1000. In the legend, the abbreviation ’CSS’ denotes our Central Server Sampling method.
Figures 2 and 3 also showcase that our algorithm outshines the current state-of-the-art across a wide array of parameters. In these experiments, we tested various graph sizes and privacy budgets, observing similar trends for both algorithms. Notably, even though the performance gap between the two algorithms appears to expand as the privacy budget increases, the difference in performance consistently remains around three orders of magnitude, irrespective of the graph size.
VII Conclusion
Our work introduces a private graph publishing mechanism that provides unbiased estimations of all graph edges while maintaining low communication costs. The effectiveness of this method stems from integrating two key components: the application of universal hashing functions to achieve uniform edge partitions and the amplification of budget efficiency through sampling within each group. This synergy leads to a significant reduction in download costs, quantifiable as a factor of $\mathcal{O}(s^{3})$, where $s$ represents the size of the groups.
We then demonstrated its utility in accurately estimating the number of triangles in a graph. Furthermore, we elaborated on the application of smooth sensitivity to these problems, ensuring that the resulting estimations remain unbiased. Our experiments have demonstrated that these methods offer substantial improvements in precision over existing state-of-the-art approaches for the counting.
Our future work is focused on enhancing the scalability of subgraph counting under local differential privacy. We have observed that the computation time required for each user in all mechanisms proposed to date is considerable. This computation time often increases for mechanisms involving multiple steps. Our aim is to explore how sampling methods can not only improve precision but also significantly boost the scalability of the counting process.
Acknowledgment
Quentin Hillebrand is supported by KAKENHI Grant 20H05965. Vorapong Suppakitpaisarn is partially supported by KAKENHI Grant 21H05845 and 23H04377. Tetsuo Shibuya is partially supported by KAKENHI Grant 20H05967, 21H05052, and 23H03345.